### v2

@@ Abstract
We computationally implement and experimentally test the behavioral predictions of a dynamic neural model of lexical meaning in the framework of Dynamic Field Theory. We demonstrate the architecture and behavior of the model using as a test case the English lexical item have, focusing on its polysemous use. In the model, have maps to a semantic space defined by two independently motivated continuous conceptual dimensions, connectedness and control asymmetry. The mapping is modeled as coupling between a neural node representing the lexical item and neural fields representing the conceptual dimensions. While lexical knowledge is modeled as a stable coupling pattern, real-time lexical meaning retrieval is modeled as the motion of neural activation patterns between transiently stable states corresponding to semantic interpretations or readings. Model simulations capture two previously reported empirical observations: (1) contextual modulation of lexical semantic interpretation, and (2) individual variation in the magnitude of this modulation. Simulations also generate a novel prediction that the by-trial relationship between sentence reading time and acceptability should be contextually modulated. An experiment combining self-paced reading and acceptability judgments replicates previous results and partially bears out the model’s novel prediction. Altogether, results support a novel perspective on lexical polysemy: that the many related meanings of a word are not categorically distinct representations; rather, they are transiently stable neural activation states that arise from the nonlinear dynamics of neural populations governing interpretation on continuous semantic dimensions.
@@ 1.1 English ‘have’
When people are presented with less frequent argument combinations, as in 1.1 below, two types of responses are reported: (a) the inanimate subject “the oak tree” leads to an inalienable possession reading, which conflicts with the fact that the object “the motorcycle” does not plausibly enter into a part-whole relation with “the oak tree”, leading to a decrease in acceptability; or (b) the referent of the subject is anthropomorphized, e.g., “the person dressed as a tree…”, which maintains an alienable possession interpretation (Zhang et al., 2018; Zhang, 2021). The facts presented so far are consistent with the hypothesis that have primarily evokes possession readings.
@@
Inclusion of an explicit locative modifier is not the only way to make an adjacency reading salient. A bias for an adjacency interpretation can also be induced by the preceding context. Specifically, when a preceding sentence evokes an adjacency reading, as in 1.1, then sentences like 1.1 are more likely to evoke an adjacency reading as well as receive higher acceptability ratings (Zhang et al., 2018, 2022), as seen in 1.1.
@@
This linguistic and behavioral pattern suggests a unified polysemy-based analysis of English have involving readings ranging from a purely coincidental spatial relation to an alienable possession relation to an inalienable possession relation (Zhang, 2021). Crucially, these readings are not disconnected. Evidence suggests that the distinction between adjacency, alienable possession, and inalienable possession is a matter of degree, not category, not only between these three reading types but also within them (e.g., Deo, 2015; Zhang, 2021; Zhang et al., 2022). In the case of coincidentality in spatial location, for example, the situation conveyed by “the tree has a motorcycle” is more coincidental than “the tree has a bench”, and that, in turn, is more coincidental than “the tree has a swing”, or “the tree has a nest”. So, with each situation we move towards less coincidentality and, interestingly, towards greater asymmetry between arguments such that the first argument is more and more perceived as “possessing” the second.
@@
We conclude then that the lexical meaning of have, while constrained, gives rise to gradable readings. As a result, such meaning is best characterized as a continuous space within which specific readings can obtain. This conclusion in turn raises the question of what the properties of such a continuous space are. We address this question below.
@@ 1.3 Real-time comprehension of have
When preceded by a context sentence designed to evoke an adjacency reading, the target sentence received higher acceptability ratings, relative to when the same target sentence was preceded by a context sentence designed to evoke a possession reading. This improvement in acceptability was attributed to a gradient shift in interpretation towards adjacency due to the influence of the context sentence.
@@ 1.4 A discrete alternative
Such a proposal is particularly appealing when only two possible readings are considered, as in the experiments described above, as well as in the experiment reported below. However, the broader range of empirical evidence reviewed above supports the continuous account: the synchronic range of have interpretation, the diachronic trajectory of change in the meaning of Marathi kade, and evidence for continuous cognitive dimensions in other domains. The continuous representational account also has two theoretical advantages over a discrete representational account: parsimony and explanatory power. Under a discrete representational account, a large number of representations would have to be posited: one for each reading. The continuous account is more parsimonious because it includes fewer representations: only two (continuous) dimensions. It is also more explanatory because it explicitly relates the different readings to each other, via direction and distance in the continuous space. Ultimately, extending our approach to the larger variety of possible interpretations of have–—as well as other cases of polysemy—–will be necessary to support the hypothesis of continuous representations for lexical meaning. A main contribution of this paper is to offer a neural processing model of lexical meaning on continuous dimensions, which is a necessary step towards extending to other interpretations and other lexical items. To situate our neural processing model, we turn now to Dynamic Field Theory (DFT: Schöner et al., 2016), a formal framework for understanding the neural activation dynamics underlying continuous cognitive dimensions.
@@ 1.5 Dynamic Field Theory (DFT)
In this subsection, we give a brief overview of DFT; a more detailed description of the relevant mathematics is given in Section 2. In DFT, features relevant for cognition are modeled as continuous parameters governed by the activity of populations of neurons. The activity of a neural population over time is described using a differential equation defining a dynamic neural field (DNF: Amari, 1977). DNFs are characterized by point attractor dynamics. This means that, at any given time, the activation pattern in a DNF is attracted to a particular state, i.e., the point attractor state. The location of the point attractor can change over time according to a variety of factors, e.g., inputs to the DNF from sensory surfaces or other DNFs. Since change within DNFs is characterized by point attractor dynamics, but the location of the point attractor changes over time, DNFs are characterized by an interplay between stability and flexibility. Stability offers resistance to the ubiquitous influence of noise, and flexibility allows rapid change under changing cognitive and environmental conditions (e.g., Kelso, 1995). Usually, when a DNF is not receiving any input, the point attractor corresponds to an inactive or resting state. When a DNF begins to be influenced by input, the point attractor might shift to an active state. In particular, the dynamics of lateral interaction within DNFs allow the formation of “peaks” of activation. Depending on the cognitive dimension being represented by the DNF, an activation peak might correspond to a movement goal, a percept, or another kind of cognitive event. Discontinuous shifts from inactive states to active states (and vice versa) exemplify the nonlinear dynamics of DNFs, whereby continuous change in one variable (e.g., input to a DNF) can result in sudden, discontinuous change in another variable (e.g., DNF activation). In the context of lexical meaning, nonlinear dynamics offer a way to reconcile intuitions of discreteness with evidence for underlying continuity.
@@ 1.6 This paper
In this study, we extend neural field dynamics to semantic dimensions underlying lexical polysemy. The model we propose implements a mapping between the lexical item have and the continuous semantic space schematized in Figures 1 and 2. In the model, peaks of neural activation correspond to semantic interpretations. One contribution of the paper is to offer a neurocognitive process-based explanation of the behavioral results described above: contextual modulation of the timecourse and outcome of lexical interpretation, and individual variation in the magnitude of this modulation. Our explanation captures the intuition that meaning is experienced as discrete (e.g., adjacency or possession), despite evidence for a continuous substrate. Such apparent discreteness, we propose, arises from nonlinear dynamics of continuous neurocognitive variables. In addition to capturing existing results, simulations from the model also generate a novel empirical prediction, which we test with an experiment combining self-paced reading and acceptability judgments. We intend our model to be general, using the English lexical item have as a test case. We propose that, while individual lexical items vary in the relevant semantic dimensions, as well as in the details of their coupling to those dimensions, the basic architecture and mechanisms of our model do not vary.
@@ 2.1 Model structure
The model consists of two dynamic neural fields (DNFs) and one dynamic node. One DNF governs interpretation on the semantic dimension of control asymmetry (CA), and the other DNF governs interpretation on the dimension of connectedness (conn). The node corresponds to the lexical item have. Activation of the node follows simple linear dynamics, given in Eq. 1.
The rate of change of activation is negatively related to current activation , defining a dynamical system with a point attractor at . represents external input to the node, e.g. from perception or intention, and represents normally distributed noise weighted by a parameter . When there is external input to the node, the node’s activation is attracted to . When there is no input, activation is attracted to 0 . is a time constant, with higher values corresponding to slower rates of evolution. In the simulations reported below, is set to , and is set to . We set the magnitude of external input depending on the condition being simulated.
Activation in each of the two DNFs unfolds according to Eq. 2.
Activation is defined for each neuron representing the relevant semantic dimension at each moment in time . The range of is set to 99, such that each neuron in the field is maximally sensitive (“tuned”) to a particular percentage of the maximum conceivable value of that dimension. For instance, when the neuron in the conn DNF is active, this corresponds to an interpretation of “ of maximum conceivable connectedness”. Activation in the DNF has a point attractor at . The resting level is assumed to be below zero for all neurons, by convention at . We set the time constant for the DNFs to . Each field input , , and is represented as a separate Gaussian distribution of the form
where controls the amplitude of the input, controls the position of the input in the field, and controls the width of the input distribution. In the simulations reported below, we set the amplitude of external input depending on the condition being simulated. represents input from the have node, whose amplitude is defined straightforwardly as a linear function of the activation of the node: . This is a simplification relative to most DFT models, where the amplitude of input from a node to a field would be a more complex (sigmoidal) function of node activation. Our motivation for eliminating the sigmoidal gating function on node-to-field input comes from work on lexical neighborhood effects on articulation, where non-selected lexical items (nodes) exert some influence on phonetic planning (DNFs) (Stern and Shaw, 2023b). This issue is largely orthogonal to the present study because there is only one node in the model. We chose to use the same node-field coupling dynamics as in Stern and Shaw (2023b) for the sake of simplicity and consistency. The same results could likely be achieved with a relatively shallow or “soft” sigmoidal gating function.
@@
represents input from each DNF to the other. Via this mechanism, patterns of activation in one DNF evoke corresponding patterns of activation in the other DNF. This mechanism implements the downward parabola pattern described in Section 1 (Figure 2). In particular, as seen in Figure 4, activation consistent with high control asymmetry evokes activation consistent with intermediate connectedness (and vice versa; note the double-sided arrows), and activation consistent with low control asymmetry evokes activation consistent with both low connectedness and high connectedness (and vice versa).
The amplitude of each input from one DNF to the other is given by
| distribution | ||
| low CA | ||
| high CA | ||
| low conn | ||
| mid conn | ||
| high conn |
@@ 2.2.1 Example simulation: alienable possession
To simulate interpretation of the sentence in 2.2, the model receives three external inputs, all with : input to the have node, input corresponding to high CA, and input corresponding to mid conn. The input to the have node corresponds to perception of the lexical item have. The other two inputs correspond to knowledge about the referents of the nouns serving as arguments of have: professor and motorcycle. We simulated model evolution under the influence of these three inputs for 90 timesteps. Figure 7 displays the results.
It can be seen that all three model components stabilize in active states shortly after the onset of the simulation. The have node is attracted to because of the external input to the node. The CA field forms an activation peak at the high end of the field because of external input centered at this location, as well as input from the conn field centered at this location (see Figure 4). The conn field forms an activation peak in the middle of the field because of external input, input from the have node, and input from the CA field. This qualitative activation state will persist until the inputs change or are removed. In this example, the model stabilizes quickly, and predictably, because the various influences on activation—external inputs, node-field coupling, and field-field coupling—reinforce each other.
@@ 2.2.2 Example simulations: contextual modulation
Next, we simulate sentences containing have with two unrelated, inanimate arguments (as in sentence 1.1) in two contexts: following an adjacency reading of have, and following a possession reading of have. These simulations are analogous to the experiment in Zhang et al. (2018), described in Section 1.3. Each simulation consists of three phases, summarized in Table 2. In phase 1, the model receives external inputs which drive interpretation of the context sentence. In the adjacency context (corresponding to sentence 1.1), these inputs correspond to low CA and low conn because of the meanings of the nouns and the locative prepositional phrase; the have node also receives input. In the possession context (corresponding to sentence 1.3), external inputs also excite the have node and the low side of the CA field, but the conn field receives input corresponding to high connectedness, i.e., inalienable possession. Again, the latter two inputs come from the meanings of the nouns which are the arguments of have. Phase 1 is the only phase which differs between conditions. In phase 2, all inputs are removed. This corresponds to the time between interpreting the context sentence and interpreting the target sentence. In phase 3, the model receives external inputs corresponding to the target sentence. These inputs are identical in both conditions: have (), low CA (), and low conn (). The amplitude of the input to the low end of the conn field is much weaker in the target sentence relative to the adjacency context sentence, because the target sentence lacks a locative prepositional phrase. Nonetheless, the two unrelated, inanimate nouns in the target sentence are assumed to induce a weak bias for a low connectedness interpretation. Phases 1 and 3 each run for 90 timesteps, which was found to be enough time for an activation peak to stabilize in both DNFs. Phase 2 runs for 20 timesteps, which was found to be enough time for both activation peaks to fall below the interaction threshold (), but not enough time for the fields to return fully to the resting level. Each simulation runs for a total of 200 timesteps.
@@
The examples in Figure 8 were selected in order to demonstrate the effect of preceding context (phase 1 evolution) on target sentence interpretation (phase 3 evolution). However, not every simulated run of the model exhibits the same effect. The presence of noise in the model introduces a stochastic influence on the location of field stabilization in each simulation. Moreover, the weak low conn input in phase 3 introduces an overall bias for low connectedness in both contexts. In order to examine the robustness of the contextual modulation effect, we simulated 1000 instances of interpretation in each of the two contexts. As seen in Figure 9, there is a bimodal distribution of interpretations in both contexts. That is, for each context, an interpretation corresponding to either adjacency or possession for the target sentence was possible. In this way, the behavior of the model can be qualitatively described as corresponding to discrete outcomes, despite the underlyingly continuous state space (continuous features, continuous time, continuous activation). The apparent discreteness of the interpretations of have is consistent with the intuitions of experiment participants (Zhang, 2021) and linguistic researchers (Myler, 2016). The discreteness arises from the structure imposed by the coupling patterns (node-field and field-field), along with the nonlinear activation dynamics which drive selection of a particular field location on each simulation. Nonetheless, there is gradient variability in the precise location of each activation peak, due to noise in each of the model components.

### v3

@@ Abstract
We computationally implement and experimentally test the behavioral predictions of a dynamic neural model of lexical meaning in the framework of Dynamic Field Theory. We demonstrate the architecture and behavior of the model using as a test case the English lexical item have, focusing on its polysemous use. In the model, have maps to a semantic space defined by two independently motivated continuous conceptual dimensions, connectedness and control asymmetry. The mapping is modeled as coupling between a neural node representing the lexical item and neural fields representing the conceptual dimensions. While lexical knowledge is modeled as a stable coupling pattern, real-time lexical meaning retrieval is modeled as the motion of neural activation patterns between transiently stable states corresponding to semantic interpretations or readings. Model simulations capture two previously reported empirical observations: (1) contextual modulation of lexical semantic interpretation, and (2) individual variation in the magnitude of this modulation. Simulations also generate a novel prediction that the by-trial relationship between sentence reading time and acceptability should be contextually modulated. An experiment combining self-paced reading and acceptability judgments replicates previous results and partially bears out the model’s novel prediction. Altogether, results support a novel perspective on lexical polysemy: that the many related meanings of a word are not categorically distinct representations; rather, they are transiently stable neural activation states that arise from the nonlinear dynamics of neural populations governing interpretation on continuous semantic dimensions. Our model offers important advantages over related models in the dynamical systems framework, as well as models based on Bayesian inference.
@@ 1.1 English ‘have’
When people are presented with less frequent argument combinations, as in 1.1 below, two types of responses are reported: (a) the inanimate subject “the oak tree” leads to an inalienable possession reading, which conflicts with the fact that the object “the motorcycle” does not plausibly enter into a part-whole relation with “the oak tree”, leading to a decrease in acceptability; or (b) the referent of the subject is anthropomorphized, e.g., “the person dressed as a tree…”, which maintains an alienable possession interpretation (Zhang et al., 2018; Zhang, 2021). The facts presented so far are consistent with the generalization that have primarily evokes possession readings.
@@
Inclusion of an explicit locative modifier is not the only way to make an adjacency reading salient, however. A bias for an adjacency interpretation can also be induced by the preceding context. Specifically, when a preceding sentence evokes an adjacency reading, as in 1.1, then sentences like 1.1 are more likely to evoke an adjacency reading as well as receive higher acceptability ratings (Zhang et al., 2018, 2022), as seen in 1.1.
@@
This linguistic and behavioral pattern suggests a unified polysemy-based analysis of English have involving readings ranging from a purely coincidental spatial relation to an alienable possession relation to an inalienable possession relation (Zhang, 2021). Crucially, these readings are not disconnected. Evidence suggests that the distinction between adjacency, alienable possession, and inalienable possession is a matter of degree, not category, not only between these three reading types but also within them (e.g., Deo, 2015; Zhang, 2021; Zhang et al., 2022). In the case of coincidentality in spatial location, for example, the situation conveyed by “the tree has a motorcycle” is more coincidental than “the tree has a bench”, and that, in turn, is more coincidental than “the tree has a swing”, or “the tree has a nest”. So, with each situation we move towards less coincidentality and, interestingly, towards greater control asymmetry between arguments such that the first argument is more and more perceived as “possessing” the second.
@@
We conclude then that the lexical meaning of have, while constrained, gives rise to gradable readings. As a result, we reason that such meaning is best characterized as a continuous space within which specific readings can obtain. This conclusion in turn raises the question of what the properties of such a continuous space are. We address this question below.
@@ 1.3 Real-time comprehension of have
When preceded by a context sentence designed to evoke an adjacency reading, the target sentence received higher acceptability ratings, relative to when the same target sentence was preceded by a context sentence designed to evoke a possession reading. This improvement in acceptability is attributed to a gradient shift in interpretation towards adjacency due to the influence of the context sentence.
@@ 1.4 A discrete alternative
Such a proposal is particularly appealing when only two possible readings are considered, as in the experiments described above, as well as in the experiment reported below. However, the broader range of empirical evidence reviewed above supports the continuous account: the synchronic range of have interpretation, the diachronic trajectory of change in the meaning of Marathi kade, and evidence for continuous cognitive dimensions in other domains. The continuous representational account also has two theoretical advantages over a discrete representational account: parsimony and explanatory power. Under a discrete representational account, a large number of representations would have to be posited: one for each reading. The continuous account is more parsimonious because it includes fewer representations: only two (continuous) dimensions. It is also more explanatory because it explicitly relates the different readings to each other, via direction and distance in the continuous space. Ultimately, extending our approach to the larger variety of possible interpretations of have–—as well as other cases of polysemy—–will be necessary to support the hypothesis of continuous representations for lexical meaning. A main contribution of this paper is to offer a neural processing model of lexical meaning on continuous cognitive dimensions, which is a necessary step towards extending to other interpretations and other lexical items. To situate our neural processing model, we turn now to Dynamic Field Theory (DFT: Schöner et al., 2016), a formal framework for understanding the neural activation dynamics underlying continuous cognitive dimensions.
@@ 1.5 Dynamic Field Theory (DFT)
In this subsection, we give a brief overview of DFT; a more detailed description of our model is given in Section 2 and A. In DFT, dimensions relevant for cognition are modeled as continuous parameters governed by the activity of populations of neurons. The activity of a neural population over time is described using a differential equation defining a dynamic neural field (DNF: Amari, 1977). DNFs are characterized by point attractor dynamics. This means that, at any given time, the activation pattern in a DNF is attracted to a particular state, i.e., the point attractor state. The location of the point attractor can change over time according to a variety of factors, e.g., inputs to the DNF from sensory surfaces or other DNFs. Since change within DNFs is characterized by point attractor dynamics, but the location of the point attractor changes over time, DNFs are characterized by an interplay between stability and flexibility. Stability offers resistance to the ubiquitous influence of noise, and flexibility allows rapid change under changing cognitive and environmental conditions (e.g., Kelso, 1995). Usually, when a DNF is not receiving any input, the point attractor corresponds to a resting state. When a DNF begins to be influenced by input, the point attractor might shift to an active state. In particular, the dynamics of lateral interaction within DNFs allow the formation of “peaks” of activation. Depending on the cognitive dimension being represented by the DNF, an activation peak might correspond to a movement goal, a percept, or another kind of cognitive event. Discontinuous shifts from inactive states to active states (and vice versa) exemplify the nonlinear dynamics of DNFs, whereby continuous change in one variable (e.g., input to a DNF) can result in sudden, discontinuous change in another variable (e.g., DNF activation). In the context of lexical meaning, nonlinear dynamics offer a way to reconcile intuitions of discreteness with evidence for underlying continuity.
@@ 1.6 This paper
In this study, we extend neural field dynamics to semantic dimensions underlying lexical polysemy. The model we propose implements a mapping between the lexical item have and the continuous semantic space schematized in Figures 1 and 2. In the model, peaks of neural activation correspond to semantic interpretations. One contribution of the paper is to offer a neurocognitive process-based explanation of the behavioral results described above: contextual modulation of the timecourse and outcome of lexical interpretation, and individual variation in the magnitude of this modulation. Our explanation captures the intuition that meaning is experienced as discrete (e.g., “adjacency” or “possession”), despite evidence for a continuous substrate. Such apparent discreteness, we propose, arises from nonlinear dynamics of continuous neurocognitive variables. In addition to capturing existing results, simulations from the model also generate a novel empirical prediction, which we test with an experiment combining self-paced reading and acceptability judgments. We intend our model to be general, using the English lexical item have as a test case. We propose that, while individual lexical items vary in the relevant semantic dimensions, as well as in the details of their coupling to those dimensions, the basic architecture and mechanisms of our model do not vary.
@@ 2.1 Model structure
The model consists of two dynamic neural fields (DNFs) and one dynamic node.888This section describes the qualitative structure and behavior of the model. Model equations with more detailed descriptions are given in A. One DNF governs interpretation on the semantic dimension of control asymmetry (CA), and the other DNF governs interpretation on the dimension of connectedness (conn). The node corresponds to the lexical item have. Activation of the node follows linear point attractor dynamics, with the position of the point attractor in activation space determined by the sum of external input to the node, e.g. from perception or intention, and random noise.
Activation in each of the two DNFs also unfolds according to point attractor dynamics. Each DNF receives inputs (e.g., from the have node, or from the other DNF, as described below) represented as Gaussian distributions with amplitude , position , and width . Under the influence of these time-varying inputs, the position of the point attractor in each DNF transitions between transiently stable resting states and transiently stable activation peaks, which correspond to semantic interpretations (see Section 1.5 and A). Crucially, we set the parameters of lateral interaction such that only a single activation peak can form at a time in each DNF, defining “selection” dynamics. The range of each DNF is set from 1 to 99, such that each neuron in each DNF is maximally sensitive (“tuned”) to a particular percentage of the maximum conceivable value of that dimension. For instance, when the neuron in the conn DNF is active, this corresponds to an interpretation of “ of maximum conceivable connectedness”.
@@
The two DNFs are also coupled to each other, such that active neurons in each DNF send input to the other DNF. Via this mechanism, patterns of activation in one DNF evoke corresponding patterns of activation in the other DNF. This mechanism implements the downward parabola pattern described in Section 1 (Figure 2). In particular, as seen in Figure 4, activation consistent with high control asymmetry evokes activation consistent with intermediate connectedness (and vice versa; note the double-sided arrows), and activation consistent with low control asymmetry evokes activation consistent with both low connectedness and high connectedness (and vice versa).
@@ 2.2.1 Example simulation: alienable possession
To simulate interpretation of the sentence in 2.2, the model receives three external inputs, all with amplitude : input to the have node, input corresponding to high CA, and input corresponding to mid conn. The input to the have node corresponds to perception of the lexical item have. The other two inputs correspond to knowledge about the referents of the nouns serving as arguments of have: professor and motorcycle. We simulated model evolution under the influence of these three inputs for 90 timesteps. Figure 5 displays the results.
It can be seen that all three model components stabilize in active states shortly after the onset of the simulation. Activation of the have node is attracted to because of the external input to the node. The CA field forms an activation peak at the high end of the field because of external input centered at this location, as well as input from the conn field centered at this location (see Figure 4). The conn field forms an activation peak in the middle of the field because of external input, input from the have node, and input from the CA field. This qualitative activation state will persist until the inputs change or are removed. In this example, the model stabilizes quickly, and predictably, because the various influences on activation—external inputs, node-field coupling, and field-field coupling—reinforce each other.
@@ 2.2.2 Example simulations: contextual modulation
Next, we simulate sentences containing have with two unrelated, inanimate arguments (as in sentence 1.1) in two contexts: following an “adjacency” reading of have, and following a “possession” reading of have.999We use the terms “adjacency” and “possession” as labels to differentiate the two context sentence types, with the understanding that, under our proposal, these labels correspond to regions within a continuous semantic space, rather than discrete representations. These simulations are analogous to the experiment in Zhang et al. (2018), described in Section 1.3. Each simulation consists of three phases, summarized in Table 1. In phase 1, the model receives external inputs which drive interpretation of the context sentence. In the adjacency context (corresponding to sentence 1.1), these inputs correspond to low CA and low conn because of the meanings of the nouns and the locative prepositional phrase; the have node also receives input. In the possession context (corresponding to sentence 1.3), external inputs also excite the have node and the low side of the CA field, but the conn field receives input corresponding to high connectedness, i.e., inalienable possession. Again, the latter two inputs come from the meanings of the nouns which are the arguments of have. Phase 1 is the only phase which differs between conditions. In phase 2, all inputs are removed. This corresponds to the time between interpreting the context sentence and interpreting the target sentence. In phase 3, the model receives external inputs corresponding to the target sentence. These inputs are identical in both conditions: have (), low CA (), and low conn (). The amplitude of the input to the low end of the conn field is much weaker in the target sentence relative to the adjacency context sentence, because the target sentence lacks a locative prepositional phrase. Nonetheless, the two unrelated, inanimate nouns in the target sentence are assumed to induce a weak bias for a low connectedness interpretation. Phases 1 and 3 each run for 90 timesteps, which was found to be enough time for an activation peak to stabilize in both DNFs. Phase 2 runs for 20 timesteps, which was found to be enough time for both activation peaks to fall below the interaction threshold (), but not enough time for the fields to return fully to the resting level. Each simulation runs for a total of 200 timesteps.
@@
The examples in Figure 6 were selected in order to demonstrate the effect of preceding context (phase 1 evolution) on target sentence interpretation (phase 3 evolution). However, not every simulated run of the model exhibits the same effect. The presence of noise in the model introduces a stochastic influence on the location of field stabilization in each simulation. Moreover, the weak low conn input in phase 3 introduces an overall bias for low connectedness in both contexts. In order to examine the robustness of the contextual modulation effect, we simulated 1000 instances of interpretation in each of the two contexts. As seen in Figure 7, there is a bimodal distribution of interpretations in both contexts. That is, for each context, an interpretation corresponding to either adjacency or possession for the target sentence was possible. In this way, the behavior of the model can be qualitatively described as corresponding to discrete outcomes, despite the underlyingly continuous state space (continuous features, continuous time, continuous activation). The apparent discreteness of the interpretations of have is consistent with the intuitions of experiment participants (Zhang, 2021) and linguistic researchers (Myler, 2016). Importantly, this apparent discreteness emerges from the structure imposed by the coupling patterns (node-field and field-field), along with the nonlinear activation dynamics which drive selection of a particular field location on each simulation. Nonetheless, there is gradient variability in the precise location of each activation peak, due to noise in each of the model components.